OpenStudy (anonymous):

The table shows the ratio of caramel corn to hot pepper corn in a bag. What is the constant of proportionality? Hot Pepper Corn Caramel Corn 5 8 10 16 15 24 3 8 1.6 0.625 The graph shows the amount of money paid when purchasing bags of caramel corn at the zoo: Write an equation to represent the relationship between the total cost (y) and the number of bags of caramel corn (x). y = 1/2x y = 2x y = 3x y = 1/3x

OpenStudy (anonymous):

@Atsie

OpenStudy (retireed):

ratio of caramel corn to hot pepper corn caramel corn / hot pepper corn 8 / 5 = ? 16 / 10 = ? 24 / 15 = ?

OpenStudy (anonymous):

for the second one

OpenStudy (retireed):

What is the slope of the line? ? RISE over RUN Change in Y divided by change in X (Y2-Y1) / (X2-X1) What is the y-intercept? What is the value of Y when X is zero?

OpenStudy (caution0):

i got a head ache just looking at this

OpenStudy (anonymous):

lol

OpenStudy (caution0):

you n eed help on the last question??

OpenStudy (anonymous):

um.. sure @caution0

OpenStudy (anonymous):

hello @Atsie @caution0

Atsie (atsie):

I really really don't understand this myself at the moment :/ Besides I have to go soon :( I can try later @ghostrider02 I'm sorry.

OpenStudy (caution0):

same,dont understand